If the marginal propensity to consume (MPC) in your classmate's nation is 3/5 and the marginal propensity to save (MPS) in your country is 1/10, which of the following must be true?​

a.

Autonomous consumption is lower in your classmate's nation than in your country.​

b.

Autonomous consumption is higher in your classmate's nation than in your country.​

c.

The spending multiplier is smaller in your classmate's nation than in your country.​

d.

Total consumption is lower in your classmate's nation than in your country.​

e.

​The spending multiplier is larger in your classmate's nation than in your country.
Answer #1

ANSWER:

The correct answer is option c that is The spending multiplier is smaller in your classmate's nation than in your country.​

spending multiplier of classmate nation = 1 / mps = 1 / (1 - mpc) = 1 / ( 1 - 3 / 5) = 1 / (2 / 5) = 2.5

spendimg multiplier in my country = 1 / mps = 1 / (1/10) = 10